OpenStudy (anonymous):

The measure of line RS is 38 and the coordinate of R is 14, what is the coordinate of S?

OpenStudy (anonymous):

|R + S| = 38 |14 + S|=38 either 14+S=38 or -(14+S)=38

OpenStudy (anonymous):

that should help

OpenStudy (anonymous):

then what is the coordinate of S?

OpenStudy (anonymous):

solve 14+S=38 and -(14+S)=38 and you will get two coordinates of S

OpenStudy (anonymous):

then the answer is -24+S? is that right?

OpenStudy (anonymous):

-24=S?

OpenStudy (anonymous):

sorry...... it should have been |R-S|=38

OpenStudy (anonymous):

which gives 14-S=38 or -(14-S)=38 S=-24 or S=52

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome